OpenStudy (anonymous):

Substitution Method 6x+7y=37 -8x+y=23

OpenStudy (campbell_st):

rewrite the 2nd equation y = 8x + 23 substitute it into the 1st equation 6x + 7(8x + 23) = 37 expand and collect like terms 6x + 56x + 161 = 27 I'll leave it to you to solve for x. When you have x substitute it into either equation to find y

OpenStudy (callisto):

* 6x + 56x + 161 = 37